OpenStudy (anonymous):

The pizza has a diameter of 12 inches. What is the closest to the area of the pizza? a. 432 square inches b. 168 square inches c. 108 square inches d. 72 square inches

OpenStudy (anonymous):

Area of a circle = pi*r^2, 2r = diameter

OpenStudy (anonymous):

2*3.14*6

OpenStudy (anonymous):

ohhh yup ur right

OpenStudy (anonymous):

2pi r

OpenStudy (anonymous):

@danny_dumond You posted the formula for the circumference, not the area.

OpenStudy (anonymous):

Cool. I keep forgetting to multiply by 2....thanks guys. C=pi*d Area of a circle=A=pi*r^2

OpenStudy (anonymous):

yes i just realized sorry

OpenStudy (anonymous):

2r = diameter r = (diameter/2) \[\pi*(\frac{diameter}{2})^{2} = \]

OpenStudy (anonymous):

Did you get 113.04?

OpenStudy (anonymous):

Yep. If you use 3 as pi you'll get 108.

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

No problemo